OpenStudy (anonymous):

Some one HELP! What method can I use to solve the equation? 2x^2 + 4x -3=0

OpenStudy (anonymous):

I no thats the reason I cant solve it!HELP

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

Quadratic formula?

OpenStudy (anonymous):

\[ x_{1/2} = \frac{-b \pm \sqrt{ b^2 - 4ac }}{2a} \]with a = 2, c = -3 and b = 4 :-)